A 5.0 kg, 22cm diameter bowling ball is rolling down an alley at 1.3 m/s. [note for a sphere: I=(2/5)MR^2]

a.)What is the ball's angular momentum

b.)What is the ball's total energy as it rolls down the alley
Answer #1

Angular momentum J = I × omega

Or J = 2/5mr2 × v /r

Or J = 2/5 mvr = 2/5 ×5×1.3×.11 0.286 kg m2 / s

B ) total energy = 1/2 mv2 + 1/2 I (omega)^2

Or E = 1/2×5×1.3^2 + 1/2×2/5×5×1.3×1.3

As v= omega ×r

So E = 4.19 Joules